OpenStudy (anonymous):

Which number is irrational? A. 16 B. 17 − − √ C. 49 − − √ D. 0.06 (repeated)

DivineSolar (divinesolar):

Ganeshie has this one :)

ganeshie8 (ganeshie8):

\[16 = \dfrac{16}{1}\] is very much rational. similarly try to interpret other options

OpenStudy (unanimoose):

A rational number can be represented as a fraction, while a decimal (such as an infinite decimal) cannot.

ganeshie8 (ganeshie8):

\[\sqrt{49} = 7 = \dfrac{7}{1}\] is ratio of two integers so it is rational too

OpenStudy (anonymous):

my bad. I meant to put A. 1/6 Not 16

ganeshie8 (ganeshie8):

Ahh okay, still 1/6 is a ratio of two integers

OpenStudy (anonymous):

So the answer is A?

OpenStudy (michele_laino):

hint: \[0.0\bar 6 = \frac{6}{{90}}\]